OpenStudy (anonymous):

hat are the center and the radius of a circle with the equation (x - 3)² + (y + 3)² = 100?

OpenStudy (trojanpoem):

I don't understand your question but center and raduis (x-m)^2 + (y-h)^2 = r^2 so r^2 = 100 r = 10 Center at (m,h) = (3,-3)